OpenStudy (anonymous):

Please help me solve (x^2y^2/x^3y^7)^-2

OpenStudy (campbell_st):

well simplify the terms inside the brackets... same base, subtract the powers.... when dividing, leave the answers in index form \[(x^{2 -3}y^{2 - 7})^{-7}\] next apply the power of a power law.. multiply the powers \[(x^a)^b = x^{a \times b}\] so the problem can be written as \[(x^{2 - 3})^{-7} \times (y^{2 - 7})^{-7}\] do the calculations of x and y... they should be both positive powers in the answer... hope it helps